# From News Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis?

### v2

@@ Abstract.
Engagement with misinformation on social media poses unprecedented threats to societal well-being, particularly during health crises when susceptibility to misinformation is heightened in a multi-topic context. This paper focuses on the COVID-19 pandemic and addresses a critical gap in understanding online engagement with multi-topic misinformation at two user levels: news sharers who share source news items on social media and post viewers who engage with online news posts. To this end, we conduct a comprehensive analysis of fact-checked source news claims related to COVID-19 and their associated posts on X, through the lens of topic diversity and conspiracy theories. We find that false news, particularly when accompanied by conspiracy theories, exhibits higher topic diversity than true news. At the news sharer level, false news has a longer lifetime and receives more posts on X than true news. Additionally, the integration of conspiracy theories is significantly associated with a longer lifetime for COVID-19 misinformation. However, topic diversity has no significant association with news sharer engagement in terms of news lifetime and the number of posts. At the post viewer level, contrary to the news sharer level, news posts characterized by heightened topic diversity receive more reposts, likes, and replies. Notably, post viewers tend to engage more with misinformation containing conspiracy narratives: false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories. Our findings suggest that news sharers and post viewers exhibit different engagement patterns on social media regarding topic diversity and conspiracy theories, offering valuable insights into designing targeted misinformation intervention strategies at both user levels.
@@ 1. Introduction
The spread of misinformation on social media platforms has emerged as one of the most significant global risks in recent years, posing far-reaching consequences for societies and individuals (Ecker et al., 2022). Health misinformation, compared to misinformation in other domains, carries unique dangers due to its potential to cause harm to human bodies and lives (Chen et al., 2022). This issue is particularly prevalent during public health crises, such as the COVID-19 pandemic, when heightened emotions like fear and anxiety, uncertainty, and increased information consumption create fertile ground for the dissemination of false or misleading information (Chuai and Zhao, 2022; Solovev and Pröllochs, 2022; Chuai et al., 2022; Chen et al., 2023; Freiling et al., 2021). Throughout the COVID-19 pandemic, a substantial amount of health misinformation regarding the virus, its origin, and potential treatments or prevention strategies spread widely on social media (Erokhin et al., 2022; Sharma et al., 2022; Teng et al., 2022). Especially, the widespread COVID-19 vaccine misinformation on social media increases beliefs in the ineffectiveness and side effects of vaccinations, contributing to individuals’ vaccination hesitancy and refusal, thereby increasing the risk of infection (Sharma et al., 2022; Pierri et al., 2022; Enders et al., 2022).
@@
Furthermore, during health crises, various aspects of society and individuals become interconnected. The discussions traverse multiple domains and demonstrate strong interconnectedness, contributing to a rich and diverse informational landscape (Miani et al., 2022; Wang et al., 2022). However, there remains a gap in understanding online engagement with misinformation within this multi-topic context during health crises. A comprehensive analysis of 13.9 million COVID-19 posts on X reveals discussions spanning more than 20 topics across the societies, encompassing areas such as politics, economy, and racism (Chandrasekaran et al., 2020). The diversity of topics within information content (i. e., multi-topic information) plays an important role in shaping perceptions and assessments of credibility through the mechanism of associative inference (Lee et al., 2023). Associative inference is an adaptive process that enables individuals to link related information and form novel connections, even in the absence of direct experience, thereby increasing individuals’ susceptibility to online misinformation (Lee et al., 2023). Moreover, the associative inference around online misinformation can evolve within a reinforcing cycle. Individuals collaborate in constructing a misleading version of reality, leading to offline actions that further reinforce this manufactured reality (Prochaska et al., 2023). The interconnected and associative multi-topic discussions during health crises underscore the importance of exploring the role of topic diversity in engagement with misinformation on social media.
Additionally, conspiracy theories (CTs) often emerge during health crises, attempting to explain the causes of significant public events as secret plots by powerful and malicious groups (Douglas and Sutton, 2018; Kou et al., 2017; Suarez-Lledo and Alvarez-Galvez, 2021). The integration of conspiracy theories within misinformation narratives serves as an effective strategy to deliberately interconnect various topics (Miani et al., 2022; Hanley et al., 2023). This deliberate approach contributes to an increase in individuals’ beliefs in misinformation by fostering illusory pattern perception, and promoting associative inference (Van Prooijen et al., 2018; Lee et al., 2023). In the early months of the COVID-19 pandemic, social media platforms were inundated with various conspiracy theories related to the origins, spread, and treatment of the virus (Erokhin et al., 2022). The propagation of misinformation with coordinated conspiracies promotes distrust in public health authorities and skepticism about the safety and effectiveness of COVID-19 vaccines (Sharma et al., 2022; Schmitz et al., 2023). While the proliferation of conspiracy theories on social media has raised significant concerns, little is known about how the incorporation of conspiracy theories in misleading narratives differentiates individuals’ engagement on social media compared to other types of misinformation (Kim and Kim, 2023). This gap calls for exploring how conspiracy theories affect engagement with misinformation during health crises.
Research questions: In this paper, we aim to provide a comprehensive understanding of online engagement with multi-topic misinformation during the early stages of the COVID-19 pandemic, focusing on topic diversity and conspiracy theories. To this end, we investigate the following Research Questions (RQs):
@@
RQ1: Does misinformation exhibit higher topic diversity compared to true information?
@@
RQ2: Does misinformation containing conspiracy theories exhibit higher topic diversity compared to misinformation without conspiracy theories?
RQ3: Does misinformation with higher topic diversity receive more engagement compared to true information?
@@
RQ4: Does misinformation containing conspiracy theories receive more engagement compared to misinformation without conspiracy theories?
@@
Of note, we examine engagement at two levels: news sharer level and post viewer level. News sharers create posts on X that link to source news items outside of the platform. Post viewers engage with news posts within the platform through the reactions of reposting, liking, or replying. The two user groups may have different motivations for engagement (Hanley et al., 2023; Teng et al., 2022; Ballard et al., 2022). Additionally, while previous studies have extensively explored post viewer engagement (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021a; Solovev and Pröllochs, 2022), research on news sharer engagement remains limited. Understanding this under-investigated area is crucial to complement our knowledge of news-sharing behavior.
@@
Methodology: We conduct our analysis based on a COVID-19 dataset consisting of fact-checked source news claims and their corresponding posts on X. These source news claims span six languages and cover the initial period of COVID-19 from January to September 2020. At news sharer level, the engagement metrics include the number of posts linking to a specific news item and the lifetime of the news item, which indicates the duration from the creation of the first post to the latest post. At post viewer level, the engagement metrics include repost count, like count, and reply count. To gain a comprehensive understanding of the topic spectrum surrounding COVID-19 discussions, we additionally use over 15 million daily sampled posts at the early stage of COVID-19 to train a Word2Vec model and collect topic keywords. Subsequently, we use the topic keywords to calculate topic diversity and identify conspiracy theories. Finally, we employ multiple regression models to investigate our research questions.
@@
Contributions: Our work reveals that elevated topic diversity has been a characteristic of COVID-19 misinformation, particularly for those accompanied by conspiracy narratives. This characteristic becomes more significant within the false news posts on X compared to their source news claims. At post viewer level, the heightened topic diversity and the integration of conspiracy theories significantly correlate with increased engagement (i. e., reposts, likes, and replies) with misinformation on X. For instance, false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories. However, at news sharer level, topic diversity and conspiracy theories have no significant association with the post count of source news items. This difference suggests that news sharers and post viewers may have different engagement patterns. Our findings offer valuable insights into understanding the engagement with multi-topic misinformation from news sharers to post viewers on social media during health crises. Additionally, by highlighting the different engagement patterns at news sharer and post viewer levels in terms of topic diversity and conspiracy theories, this research informs the development of targeted interventions and strategies at both user levels to mitigate engagement with false information and foster a more resilient society.
@@ 2.1. Misinformation on Social Media
Social media platforms have transformed information exchange, enabling public users to create and engage with online content globally and instantaneously, albeit without the quality control mechanisms inherent in traditional media. This has led to a significant rise in misinformation. Particularly, the prevalence of misinformation on social media has posed far-reaching consequences across many critical societal domains ranging from politics to public health, becoming a concerning global issue in recent years (Ecker et al., 2022; Pierri et al., 2022; Enders et al., 2022; Green et al., 2022). Consequently, social media platforms are under increasing pressure to implement measures to reduce the spread of misinformation. Understanding how users engage with misinformation and contribute to its viral spread on social media can help social media platforms to design targeted interventions. Prior research has shown that online misinformation spreads farther, faster, deeper, and more broadly compared to true information, as evidenced by the examination of online posts fact-checked by third-party organizations (Vosoughi et al., 2018). Nevertheless, another study, based on community fact-checked posts on X, indicates that misinformation receives fewer reposts compared to true information (Drolsbach and Pröllochs, 2023). This contradiction highlights that not every piece of misinformation spreads faster than true information. Instead, there are likely crucial factors behind the veracity of information that make misinformation unique and contribute to its online diffusion.
Several works have focused on understanding what characterizes misinformation and contributes to its viral spread. For instance, misinformation is more likely to incorporate negative emotions, such as anger, which have been identified as a main driver for information diffusion (Chuai and Zhao, 2022; Chuai et al., 2022; Horner et al., 2021; Robertson et al., 2023b). Additionally, there is a partisan asymmetry that misinformation tends to be right-leaning and promote conservative positions (Chuai et al., 2023). Social media users are more likely to believe and share information that is congruent to their political ideologies (Garrett and Bond, 2021; Rao et al., 2022; Robertson et al., 2023a; Osmundsen et al., 2021; Jenke, 2023). Therefore, right-leaning users are more likely to share misinformation on social media, compared to left-leaning users (Nikolov et al., 2021). More importantly, uncertain situations, such as emergency public events or crises, create fertile ground for online misinformation to thrive. During such times, people are eager to know more information about the situations and thus more likely to believe and share related misinformation, which can destroy the trust in the control policies and cause detrimental harm to societies and individuals.
@@ 2.2. Misinformation During Public Events/Crises
During public events or crises, characterized by heightened emotions, uncertainty, and a surge in information consumption, falsehoods that cater to specific demands and offer related information to contextualize the situation become rampant and easily spread (Chuai and Zhao, 2022). The preservation of democracy is crucial during political elections. However, democracy can be undermined by the widespread dissemination of dedicated online misinformation aimed at pursuing specific political agendas. It has been found that people are exposed to misinformation more often during elections than usual (Allcott and Gentzkow, 2017). Election-related misinformation is often polarized, which can increase people’s susceptibility and affect their voting behaviors (Rathje et al., 2021; Nikolov et al., 2021; Recuero et al., 2020). Additionally, climate change poses challenges at multiple levels, with any mitigation efforts carrying significant economic and political implications (Lewandowsky, 2021). Consequently, the debate surrounding global climate change has persisted for years and has become increasingly polarized on social media (Fischer et al., 2022; Farrell, 2019; Falkenberg et al., 2022). This polarized environment fosters climate change denial and misinformation, further complicating efforts to address and control climate change (Lewandowsky, 2021; Falkenberg et al., 2022).
The spread of misinformation during health crises is particularly severe, as it can directly harm human health and even cost lives (Chen et al., 2022). Throughout the COVID-19 pandemic, a significant volume of misinformation regarding a wide variety of COVID-19 related topics has circulated widely on social media platforms. For example, the widespread COVID-19 vaccine misinformation reduces users’ beliefs in its effectiveness, impeding the fight against the spread of COVID-19 (Enders et al., 2022; Loomba et al., 2021; Pierri et al., 2022). A global survey has revealed that a substantial part of people consider COVID-19 misinformation as highly reliable (Roozenbeek et al., 2020). Understanding how misinformation spreads and engages users during such public events and crises is crucial for developing effective intervention strategies for emergency management.

### v3

# From News Source Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis

@@ Abstract.
Online engagement with misinformation threatens societal well-being, particularly during health crises when susceptibility to misinformation is heightened in a multi-topic context. Here, we focus on the COVID-19 pandemic and address a critical gap in understanding engagement with multi-topic misinformation on social media at two user levels: news source sharers (who post news items) and post viewers (who engage with news posts). To this end, we analyze fact-checked source news items and their associated posts on X through the lens of topic diversity and conspiracy theories. We find that false news, especially those containing conspiracy theories, exhibits higher topic diversity than true news. At news source sharer level, false news has a longer lifetime and receives more posts on X than true news, with conspiracy theories further extending its longevity. However, topic diversity does not significantly influence news source sharers’ engagement. At post viewer level, contrary to news source sharer level, posts characterized by heightened topic diversity receive more reposts, likes, and replies. Notably, post viewers tend to engage more with misinformation containing conspiracy narratives: false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to those without conspiracy theories. Our findings suggest that news source sharers and post viewers exhibit distinct engagement patterns on X, offering valuable insights into refining misinformation interventions at these two user levels.
@@ 1. Introduction
The spread of misinformation – i. e., general false or misleading information created either intentionally or unintentionally – on social media platforms has emerged as one of the most significant global risks in recent years, posing far-reaching consequences for societies and individuals (Ecker et al., 2022; Chuai et al., 2024b). Health misinformation, compared to misinformation in other domains, carries unique dangers due to its potential to cause harm to human bodies and lives (Chen et al., 2022). This issue is particularly concerning during public health crises, such as the COVID-19 pandemic, when heightened emotions, uncertainty, and increased information consumption create fertile ground for the dissemination of false or misleading information (Chuai and Zhao, 2022; Solovev and Pröllochs, 2022; Chuai et al., 2022; Chen et al., 2023; Freiling et al., 2023). Additionally, as a distinct type of misinformation, conspiracy theories are deliberately fabricated and attempt to explain the causes of significant public events as secret plots by powerful and malicious groups (Douglas and Sutton, 2018; Kou et al., 2017; Suarez-Lledo and Alvarez-Galvez, 2021; Hanley et al., 2023). Throughout the COVID-19 pandemic, misinformation and conspiracy theories regarding the virus, its origin, and potential treatments or prevention strategies have spread widely on social media, undermining pandemic control efforts (Erokhin et al., 2022; Sharma et al., 2022b; Teng et al., 2022). For instance, widespread anti-vaccine conspiracy theories on social media have strengthened beliefs in the ineffectiveness and side effects of vaccinations, contributing to individuals’ vaccination hesitancy and refusal, thereby increasing the risk of COVID-19 infection (Sharma et al., 2022b; Pierri et al., 2022; Enders et al., 2022).
Notably, during health crises, various aspects of society become interconnected. The discussions traverse multiple domains and demonstrate strong interconnectedness, contributing to a rich and diverse informational landscape (Miani et al., 2022; Wang et al., 2022). For example, a comprehensive analysis of 13.9 million COVID-19 posts on X (formerly Twitter) reveals discussions spanning more than 20 topics across society, encompassing areas such as politics, economy, and racism (Chandrasekaran et al., 2020). Yet, an understanding of how users engage with misinformation within the multi-topic context during health crises is still in its infancy. Vosoughi et al. (2018) examined a large-scale of true and false news posts distributed on X from 2006 to 2017, finding that false news posts are more novel and attract more replies expressing surprise compared to true news posts. The novelty in their study was measured by the topic similarity between the true (false) news post a user reposted and the previous posts to which the user had been exposed before the repost. However, other research shows that social media users often have specific topic preferences and tend to consume similar content, especially within groups of like-minded individuals, i. e., echo chambers (Cinelli et al., 2021). Additionally, social media recommendation algorithms frequently feed users content based on their view histories and increase echo chambers, further reinforcing users’ preferences (Cinus et al., 2022; Engel et al., 2023). Reconciling these seemingly conflicting findings, one possible explanation is that false news may strategically connect users’ interests with other topics to pursue (hidden) agendas. Particularly, in the context of health crises marked by high uncertainty, diverse topics are more interconnected within discussions around crises (Chandrasekaran et al., 2020). Therefore, an analysis of topic interconnectedness in true and false news items during health crises is warranted.
@@
Previous research has shown that the diversity of topics within information content (i. e., multi-topic information) plays an important role in shaping perceptions and assessments of credibility through the mechanism of associative inference (Lee et al., 2023). Associative inference is an adaptive process that enables individuals to link related information and form novel connections, even in the absence of direct experience, thereby increasing individuals’ susceptibility to online misinformation (Lee et al., 2023). Moreover, the associative inference around online misinformation can evolve within a reinforcing cycle. Individuals collaboratively construct a misleading version of reality and amplify alleged evidence. This facilitates subsequent actions that, in turn, further reinforce the manufactured reality (Prochaska et al., 2023). Therefore, the interconnected and associative multi-topic discussions during health crises underscore the potential role of topic diversity in driving engagement with misinformation on social media. Based on this rationale, we address the following research questions to explore topic diversity in misinformation (RQ1.1) and examine whether it shapes user engagement (RQ1.2):
@@
RQ1.1: Does misinformation exhibit higher topic diversity compared to true information?
RQ1.2: Is topic diversity positively associated with user engagement, and is this association stronger for misinformation than for true information?
Moreover, the integration of conspiracy theories within misinformation narratives serves as an effective strategy to deliberately interconnect various topics (Miani et al., 2022; Hanley et al., 2023). This deliberate approach contributes to an increase in individuals’ beliefs in misinformation by fostering illusory pattern perception and promoting associative inference (Van Prooijen et al., 2018; Lee et al., 2023). In the early months of the COVID-19 pandemic, social media platforms were inundated with various conspiracy theories related to the origins, spread, and treatment of the virus (Erokhin et al., 2022). The propagation of misinformation with coordinated conspiracies has promoted distrust in public health authorities and skepticism about the safety and effectiveness of COVID-19 vaccines (Sharma et al., 2022b; Schmitz et al., 2023). While the proliferation of conspiracy theories on social media has raised significant concerns, little is known about how the incorporation of conspiracy theories in misleading narratives differentiates individuals’ engagement on social media compared to other types of misinformation (Kim and Kim, 2023). Given this, we further investigate the following research questions to analyze whether misinformation containing conspiracy theories exhibits higher topic diversity (RQ2.1) and receives more engagement (RQ2.2) compared to misinformation without conspiracy theories:
@@
RQ2.1: Does misinformation containing conspiracy theories exhibit higher topic diversity compared to misinformation without conspiracy theories?
@@
RQ2.2: Does misinformation containing conspiracy theories receive more engagement compared to misinformation without conspiracy theories?
@@
To address our research questions, we perform a comprehensive analysis of online engagement with multi-topic misinformation during the early stages of the COVID-19 pandemic, focusing on topic diversity and conspiracy theories. It is worth noting that we examine engagement at two levels: news source sharer level and post viewer level. News source sharers create posts on X that link to source news items, while post viewers encounter these news posts within the platform and then engage by reposting, liking, or replying. The two user groups may have different motivations for engagement (Hanley et al., 2023; Teng et al., 2022; Ballard et al., 2022). Additionally, while previous studies have extensively explored post viewer engagement on social media (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021a; Solovev and Pröllochs, 2022), research on news source sharer engagement remains limited. Understanding this under-investigated area is crucial to complement our knowledge of news-sharing behavior.
@@
Methodology: We conduct our analysis based on a COVID-19 dataset consisting of fact-checked source news claims and their corresponding posts on X. These source news claims span six languages and cover the initial period of COVID-19 from January to September 2020. At news source sharer level, the engagement metrics include the number of posts linking to a specific news item and the lifetime of the news item from the creation of its first post to its latest post. At post viewer level, the engagement metrics include repost count, like count, and reply count. To gain a comprehensive understanding of the topic spectrum surrounding COVID-19 discussions, we additionally use over 15 million daily sampled posts at the early stage of COVID-19 to train a Word2Vec model and collect topic keywords. Subsequently, we use the topic keywords to calculate topic diversity and identify conspiracy theories. Finally, we employ multiple regression models to investigate our research questions.
@@
Contributions: Our work reveals that elevated topic diversity has been a characteristic of COVID-19 misinformation, particularly for those accompanied by conspiracy narratives. This characteristic becomes more significant within false news posts on X compared to their source news claims. At post viewer level, heightened topic diversity and the integration of conspiracy theories significantly correlate with increased engagement (i. e., reposts, likes, and replies) with misinformation on X. For instance, false news posts that contain conspiracy theories, on average, receive 40.8% more reposts, 45.2% more likes, and 44.1% more replies compared to false news posts without conspiracy theories. However, at news source sharer level, topic diversity and conspiracy theories have no significant associations with the number of posts that source news items receive. This difference suggests that news source sharers and post viewers have distinct engagement patterns. Our findings offer valuable insights into understanding engagement with multi-topic misinformation from news source sharers to post viewers on social media during health crises. Additionally, by highlighting the different engagement patterns at news source sharer and post viewer levels in terms of topic diversity and conspiracy theories, this research informs the development of targeted interventions and strategies at both user levels to mitigate engagement with false information and foster a more resilient society.
@@ 2.1. Misinformation on Social Media
Social media platforms have transformed information exchange, enabling public users to create and engage with online content globally and instantaneously. However, the absence of quality control mechanisms inherent in traditional media has led to a significant rise in misinformation on social media. Particularly, the spread of misinformation on social media has posed far-reaching consequences across many critical societal domains ranging from politics to public health, thus becoming a concerning global issue in recent years (Ecker et al., 2022; Pierri et al., 2022; Enders et al., 2022; Green et al., 2022). Consequently, social media platforms are under increasing pressure to implement effective measures to reduce the spread of misinformation. Understanding how users engage with misinformation and contribute to its viral spread on social media can help social media platforms design targeted interventions. Prior research has shown that online misinformation spreads farther, faster, deeper, and more broadly compared to true information, as evidenced by the examination of online posts fact-checked by third-party organizations (Vosoughi et al., 2018). Nevertheless, another study, based on community fact-checked posts on X, indicates that misinformation receives fewer reposts compared to true information (Drolsbach and Pröllochs, 2023). This contradiction highlights that not every piece of misinformation spreads faster than true information. Instead, there are likely crucial factors behind the veracity of information that make misinformation unique and contribute to its online diffusion.
Several works have focused on understanding what characterizes misinformation and contributes to its viral spread. For instance, misinformation is more likely to incorporate negative emotions, such as anger, which has been identified as an important driver for information diffusion (Chuai and Zhao, 2022; Chuai et al., 2022; Horner et al., 2021; Robertson et al., 2023b). Additionally, there is a partisan asymmetry that misinformation tends to be right-leaning and promote conservative positions (Chuai et al., 2025). Social media users are more likely to believe and share information that is congruent to their political ideologies (Garrett and Bond, 2021; Rao et al., 2022; Robertson et al., 2023a; Osmundsen et al., 2021; Jenke, 2023). Therefore, right-leaning users are more likely to share misinformation on social media, compared to left-leaning users (Nikolov et al., 2021). More importantly, uncertain situations, such as public emergencies or crises, create fertile ground for online misinformation to thrive. During such times, people are eager to know more information about the situation and are thus more likely to believe and share related misinformation, which can undermine trust in control policies and cause detrimental harm to societies and individuals.
@@ 2.2. Misinformation During Public Events/Crises
During public events or crises, characterized by heightened emotions, uncertainty, and a surge in information consumption, falsehoods that cater to specific demands and offer related information to contextualize the situation become rampant and easily spread (Chuai and Zhao, 2022). The preservation of democracy is crucial during political elections. However, democracy can be undermined by the widespread dissemination of dedicated online misinformation aimed at pursuing specific political agendas. It has been found that people are exposed to misinformation more often during elections than usual (Allcott and Gentzkow, 2017; Chuai et al., 2025). Election-related misinformation is often polarized, which can increase people’s susceptibility and affect their voting behaviors (Rathje et al., 2021; Nikolov et al., 2021; Recuero et al., 2020). Additionally, climate change poses challenges at multiple levels, with any mitigation efforts carrying significant economic and political implications (Lewandowsky, 2021). Consequently, the debate surrounding global climate change has persisted for years and has become increasingly polarized on social media (Fischer et al., 2022; Farrell, 2019; Falkenberg et al., 2022). This polarized environment fosters climate change denial and misinformation, further complicating efforts to address and control climate change (Lewandowsky, 2021; Falkenberg et al., 2022).
The spread of misinformation during health crises is particularly severe, as it can directly harm human health and even cost lives (Chen et al., 2022). Throughout the COVID-19 pandemic, misinformation on a wide variety of COVID-19 related topics has circulated widely on social media platforms. For example, the widespread misinformation on the COVID-19 vaccine reduces users’ beliefs in its effectiveness, impeding the fight against the spread of COVID-19 (Enders et al., 2022; Loomba et al., 2021; Pierri et al., 2022). A global survey has revealed that a substantial part of people consider COVID-19 misinformation as highly reliable (Roozenbeek et al., 2020). Understanding how misinformation spreads and attracts users during such public events and crises is crucial for developing effective intervention strategies for emergency management.